OpenStudy (anonymous):

If sin theta=(-3/5) and theta terminates in the fourth quadrant, find the exact value of tan 2theta

OpenStudy (anonymous):

1. If the angle is in the Fourth Quadrant, then we know that... sin(θ) < 0 and cos(θ) > 0. Then... cos(θ) = 4/5 sin(θ) = -3/5 tan(θ) = sin(θ)/cos(θ) = -3/4 Hence... tan(2θ) = 2tan(θ)/(1 - tan²(θ)) = 2(-3/4)/(1 - (-3/4)²) = -24/7

OpenStudy (anonymous):

if that is right...I believe it is...we just learned this

OpenStudy (anonymous):

Yeah it is, thanks!

OpenStudy (anonymous):

No Problem :) glad i could help